Subject: Minutes of 1998 July 13 APO 3.5m User's Committee meeting

From: strauss@astro.Princeton.EDU

Submitted: Thu, 16 Jul 1998 09:09:42 -0400 (EDT)

Message number: 286 (previous: 285, next: 287 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************July 13, 1998**********************

Attending: Michael Strauss, Bruce Gillespie, Ed Turner, Jeff Brown,
Chris Stubbs, Alan Uomoto, Rene Walterbos,  Ed Kibblewhite

Agenda/table of contents:
Shutdown Status/Progress

	- Echelle delays
	- SPIcam filter wheel scheme
	- DIS electronics
	- Visitor instrument lessons

Operations Issues
	- REMARK vs. command line
	- OS observing duties
	- Automated Web schedules

Organizational Matters
	- Annual Community Meeting
	- Computer Committee

Appendix A: List of shutdown activities. 
Appendix B: Plan for SPICAM filter wheels.
Appendix C: Draft operational requirements for visiting instrument teams

  Gillespie: There are roughly two dozen items on the list of tasks
for the shutdown.  Appendix A below gives the full list and their
present status.  Half of the tasks have been started.  Most of the
relevant materials are in hand; we should be able to get through the
full list by the end of the shutdown.
  Most of the tasks are maintenance items, with some notable
exceptions; for example, a conical baffle system will be put in at the
2nd Nasmyth port; this should cut down on scattered light quite a bit.
The baffling will not obscure the guider.

  Also, Stubbs et al paid a recent visit to the WIYN telesope to study
their successful baffling; we are thinking about simply copying their system. 
  Note: monsoons came right on time; the weather has been bad for
the entire shutdown thus far (so no good observing has been lost to
the shutdown).  Now it simply needs to clear up at the end of the shutdown...

   The echelle has seen delays due to collimation problems, or perhaps
a problem with the figure of the Schmidt corrector plate.  The
optimistic estimate is for a delivery in September '98; no general
usage before 1st quarter 1999.
  Of course, if the corrector plate needs refiguring, then the delay
will be quite a bit longer...
  Gillespie, Turner, and Stubbs are planning a trip to Chicago to
assess the situation. 
  There is much interest, and anxiety in the community in using the
echelle, and a general need to have the most realistic schedule
possible for when it might be available. 

   SPICAM: There is a plan in place to deal with the large number of
filters being used with SPICAM, for an estimated $25,000; a series of
4 wheels, to cut down on handling of filters (to protect the filters
themselves), together with some critical spares.  The plan is appended
below in Appendix B.  Notice that this discussion is decoupled from
that of Stubbs' wide-field imaging camera, which will not have the
option of rapid filter changes (note that SPICAM will stay at the
site, even after the wide-field camera arrives).  The changing of
filter wheels is of order a five-minute overhead.  As the SDSS filters
are the most delicate, we should make it a hard rule that they stay in
their wheel no matter what.  Users should comment if they find this
too restrictive for their science..
  There is a need a proper storage box for the filter wheels; Stubbs
and APO operations are looking at it. 

   DIS electronics: Jim Gunn and Connie Rockosi have had a quick look;
with a replacement of some of the electronics with spares, the high
readnoise on the red side may be fixed.  A/D converter has been
replaced just over the last few days; we should find out soon whether
this was successful. 

  There were 2 successful visitor instruments over the last quarter;
based on these experiences, Gillespie et al have drafted a list of
operational guidelines for visitor instruments (see Appendix C
below), to make everybody's life easier.  Most important is the first
item: a single point of contact and responsibility to act as liaison
between instrument team and ARC staff.  Much discussion of how this
would work: Stubbs suggests drafting a brief document describing to a
PI wanting to bring in a visitor instrument what they are expected to

Operations Issue:
  There are some people who observe without REMARK, sending commands
directly to the MC.  There have been some problems with this, mostly
due to miscommunication between observer and observing specialist
(both sitting there, waiting for the other to do something).  Some are
due to strange glitches that are impossible to reproduce later.  The bottom
line: observers should know that there are nasty paths that one can go
down when communicating to the MC directly, and there is no guarantee
that all will work. With planned upgrades to the remote observing
interface, the need to use alternative control paths to the telescope
and instruments should be mitigated.

  Observing specialist observing duties: There have been times when
the observing specialists have been asked to take over the observing
itself.  One possible problem is that some observers are less
knowledgeable about the operation of the telescope than others.
Perhaps this is because there is not a clear description of what a new
observer must actually learn when first at the telescope.  Another
driver is that it really can be more efficient to work closely with
the observing specialist than for the observer to do everything him or
herself (for example, putting a faint object on the slit).  Much
discussion of how to facilitate and make more efficient the remote
observing capabilities.

  Automated Web Schedules: Craig Loomis has written code to turn the
ascii schedule into updateable hyperlinked text. This will be put on
the Web soon, eventually to replace the current ascii
schedule. [Editor's note: This has now happened, see apo35-general
#285, and  Take a look and
send comments via the e-mail link provided.] 

Organizational matters:
   Last year, we had a highly successful meeting of the APO user's
community, and decided to make this an annual affair.  Nothing has
been planned thus far.   Various possibilities are discussed; the
tentative plan is for Stubbs to host the meeting in late September or
mid October.  Watch this space for more details. 

Computer Committee: 
  Ed Turner is calling for nominations for people from the various
institutions to sit on this committee.  They would write
specifications for the software update for the operating systems.   UC
representatives should all nominate participants on this committee
from their institutions. 

  A seeing monitor has been prototyped on the site; there is a possibility of
sharing this with the SDSS folks. 

  Last month's minutes are approved.

  No meeting scheduled for August.  Next meeting, Tuesday, September
8, 12:30 PM 

Appendix A: Plans for July shutdown

		July 1998 Engineering Shutdown Activities

[Status as of 7/13/98, per Jon Davis]

Mount Alt Axis and Rot Control Cabinets on left rear wall of instrument
pit.  Shorten and reroute cables, remount components for easy removal
(rivnuts), cleanup of wires and cables, check grounds, route in cable
ducting, make room for additional components for telescope monitoring
system and controller for tertiary rotation.
START:  7/6/98; FINISH:  7/15/98
STATUS:  Started, progress OK

Thorough cleaning and relubrication of az axis bearing.
START:  7/8/98; FINISH:  7/10/98
STATUS:  Not started, materials acquired

Service pumps for PMSS.
START:  7/10/98; FINISH:  7/10/98
STATUS:  Not started, materials acquired

Relocate pumping station for PMSS from current location to instrument rack.
START:  7/13/98; FINISH:  7/17/98
STATUS:  Not started, materials acquired

Install monitor system cables for PMSS and retune using monitor system.
START:  7/20/98; FINISH:  7/21/98
STATUS:  Not started, materials acquired

Telescope mechanical:
Rotator:  Remove to access encoder, inspect/replace encoder bearings, check
large bearing starting torque, check rot motor.
START:  7/15/98; FINISH:  7/22/98
STATUS:  Not started

Alt:  Inspect/repair drive and encoder.  Modify magna sensor pick-up head
to "follow" disc RO.
START:  7/17/98; FINISH:  7/23/98
STATUS:  Not started

Az:  Inspect/repair drives, idlers, and encoder if time permits.
START:  7/??/98; FINISH:  7/??/98
STATUS:  Not started

Modify and reattach floor skirt around telescope.  Still contacts on right
rear corner, with rear section loose.  Tabs need to be removed for ease of
access to az drives.
START:  7/6/98; FINISH:  7/17/98
STATUS:  50% completed

Rotator Limit Switch installation and integration (switch mount design and
fabrication required).
START:  7/  /98; FINISH:  7/  /98
STATUS:  Not started

Rotator code mods for limit switch and magnasensor (encoder scale algorithm
START:  7/  /98; FINISH:  7/  /98
STATUS:  Not started

Instrument pit louver installation and integration.
START:  7/6/98; FINISH:  7/17/98
STATUS:  Not started

Remake cable guide bracket for rotator (bracket design and fabrication
START:  7/6/98; FINISH:  7/10/98
STATUS:  10% completed

Temperature Sensor Cabling work.
START:  7/??/98; FINISH:  7/??/98

Brackets for Lamps, and lamp relocation.
START:  7/6/98; FINISH:  7/10/98
STATUS:  20% completed

NA2 Baffle Installation (contingent upon completion of baffles).
START:  7/??/98; FINISH:  7/??/98
STATUS:  Baffles fabricated, installation starts this week

Support for Echelle installation???
START:  7/??/98; FINISH:  7/??/98
STATUS:  Cancelled

Mount Guider Electronics on stand (attached to telescope) and clean up cabling.
START:  7/  /98; FINISH:  7/  /98
STATUS:  10% completed

"Moat" maintenance (should be done prior to bearing cleaning and
relubrication).  Drain and filter antifreeze.  Rinse out moat.  Refill with
filtered antifreeze and top off with fresh antifreeze.  Check on disposal
requirements and have container on hand for disposal.
START:  7/6/98; FINISH:  7/7/98
STATUS:  100% completed

Concrete pier PVC duct sealing.  Cut PVC flush with concrete and install
seals.  Modify and install seal for cable.
START:  7/6/98; FINISH:  7/6/98
STATUS:  100% completed

Enclosure rotation inspections, cleaning, and checkouts.  Clean thoroughly
rail and wheels, check cam followers, bolt tightness, oil in gear drive
(change), and add window to controller enclosure (need to order).
START:  7/8/98; FINISH:  7/8/98
STATUS:  75% completed

Repair mirror covers where the rivets have broken.  New rivets have been
acquired and painted.
START:  7/  /98; FINISH  7/  /98
STATUS:  100% completed

Mount optical bench at NA1 for tests of echelle collimation and focal position
START:  7/  /98; FINISH  7/  /98
STATUS:  Bench mounted and beam collimated, waiting sky tests for focal

Appendix B: Plan for SPICAM filter wheel, from Bruce Gillespie

Chris and I devised a plan for SPIcam filters that we wish to open for
discussion.  Implementation will follow, pending identification of some

This plan has several advantages.  It minimizes filter handling (which is
becoming more and more a time-sink and risk to the filters), provides for
spares of critical filters, and makes for easy and fast filter
configuration changes.

The proposal calls for four filter wheels.

	Sloan wheel:  contains 5 SDSS filters, and is never messed with.

	Johnson wheel:  contains Johnson UBVRI, and is never messed with.

	Mix-and-match wheel:  can be configured with 24 hours advance notice
	from "stock" filters (but not those from the Sloan or Johnson wheels).

	Stubbs wheel:  reserved for Stubbs

There would be a storage facility for these wheels near the telescope to
facilitate cleanliness and wheel switching at night.

We would have to buy a new set of SDSS and Johnson filters to do this, plus
an additional wheel.  I am looking into the price of filters, but I suspect
that we would spend close to $20k +/- to implement this plan in its
entirity.  I thought that there might be some extra SDSS filters floating
around the SDSS world, but it turns out that they are 2 x 2 inch, not 3 x 3.

Appendix C:  Draft operational requirements for visiting instrument teams
		       Bruce Gillespie

	o a person from an ARC institution needs to be single point of
contact before, during, and after runs.  This person needs to be on site
before the run for pre-run planning, and during the run(s) to provide
liaison support between the instrument team and APO operations.

	o first-time APO runs should be split into two, with a 4-week break
between runs.

	o proper focal plane position and alignment requirements must be
provided to visiting instrument team prior to run.

	o the importance of the instrument block and the procedure for
determining it must be communicated to the visiting instrument team prior
to the run.

	o expectations on the level and scope of user support that APO
provides to instrument teams must be established before the run.

	o runs should be scheduled for first half nights because the teams
tend not to sleep during the day and will work all day and all night if

	o runs should not start nor stop on weekends.

	o visiting instruments should not be commissioned on this telescope
unless there are compelling reasons, e.g., no other available telescope.

	o a "trip report" and/or sample data needs to be provided to APO
after a run, to help promote visiting instrument programs in general, and
to furnish us with user feedback.

	o must require that visiting instrument team, through the ARC point
of contact, provide APO with all answers to technical and logistical
questions at least a month before run starts.

	o visiting instrument team needs to be limited to no more than
three people, including the ARC liaison, unless compelling reasons

	o visiting runs should be at least three half nights, and longer
for first time at telescope.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 286 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to